a junior College awarded 26 varsity letters in crew, 15 in swimming, and 16 in soccer. If awards went to 46 students and only 2 lettered in all sports how many students lettered in 2 of the 3 sports?: a junior College awarded 26 varsity letters in crew, 15 in swimming, and 16 in soccer. If awards went to 46 students and only 2 lettered in all sports how many students lettered in 2 of the 3 sports?
Answer by ankor@dixie-net.com(4660)

a junior College awarded 26 varsity letters in crew, 15 in swimming, and 16 in soccer. If awards went to 46 students and only 2 lettered in all sports how many students lettered in 2 of the 3 sports?
:
Eliminate the 2 that lettered in all 3 sport
46 - 2 = 44 students
:
26 - 2 = 24 crew
15 - 2 = 13 swim
16 - 2 = 14 soccer
-----------
total = 51 awards given to 44 students
:
51 - 44 = 7 students lettered in two sports
:
Check 7(2) + 37(1) = 51